OpenStudy (anonymous):

What is the slope of a line that is perpendicular to the line whose equation is 2y = 3x - 1?

OpenStudy (anonymous):

y= 3/2x -1/2 (I just divided 2 on both sides to get y by itself) The slope is 3/2 So a slope of a line perpendicular to would start like y= -2/3 which is the opposite

OpenStudy (anonymous):

Did they give u more info?

OpenStudy (anonymous):

Its -3 right?

OpenStudy (anonymous):

what is -3?

OpenStudy (anonymous):

the answer

OpenStudy (anonymous):

the answer would be y=-2/3 + ????

OpenStudy (anonymous):

Did they give you more infO?

OpenStudy (anonymous):

No that equation is all they gave me

OpenStudy (anonymous):

It doesnt say find a line that goes thru (?,?) ?

OpenStudy (anonymous):

No thats it

OpenStudy (anonymous):

Is it multiple choice?

OpenStudy (anonymous):

-3 -1/3 -2/3

OpenStudy (anonymous):

WOW, all u need to do is find the slope? I gave u the answer already.

OpenStudy (anonymous):

-2/3 haha

OpenStudy (anonymous):

yeah, lol.. i didn't know what u wanted me to find, but u get it? or u want me to elaborate?

OpenStudy (anonymous):

uh no u dont need to

OpenStudy (anonymous):

but thanks!